OpenStudy (anonymous):

1.a 100 kg motorcycle carrying a 70 kg rider comes to a stop in 40m from a speed of 50km/h when it brakes applied.find the force exerted by the brakes and the force experienced by the rider. 2.a 500g ball lying on the ground is kicked with a force of 250 N.If the kick lasted 0.020s, the ball flew of with a speed of: a.)0.01 m/s b.)2.5 m/s c.)0.1m/s d). 10 m/s 3.the coefficient of static and sliding friction for wood on wood are respectively 0.50 & 0.30. If a 100lb wooden box is pushed across a horizontal wooden floor with just enough force of static friction,its acceleration is?

OpenStudy (anonymous):

these are the last ones I need an answer to...-.- my brain wont function well anymore after 62 questions in physics..pls help thanks

OpenStudy (anonymous):

One by one okay?

OpenStudy (anonymous):

ok...sorry for posting more than one question....O.O....sorry >.<

OpenStudy (anonymous):

1) first convert 50km/h into m/s. You must calculate the acceleration experienced by the rider and/or the bike (they are equal) Use \[\Large v_{f}^2=v_{i}^2+2as\] to calculate the acceleration. Then the force experienced by the rider is \[\Large F_{rider}=m_{rider}a\] Force experienced by the bike is \[\Large F_{bike}=m_{bike}a\]

OpenStudy (anonymous):

Confirm that you understand it before I move on to the next question.

OpenStudy (anonymous):

so thats 13.89 m/s 0 m/s= Vi+ 2 (a)(13.89m/s) are my values correct?

OpenStudy (anonymous):

welp they are squared-.-" so: (0m/s)^2 = (Vi)^2 + 2(a)(13.89m/s)

OpenStudy (anonymous):

No, vi=initial velocity=13.89m/s s=distance traveled=40m.

OpenStudy (anonymous):

woups...(0m.s)^2 = (13.89m/s)^2 + 2(a)(40m)

OpenStudy (anonymous):

correct. :)

OpenStudy (anonymous):

sorry for my thinking its already 4 am here and im still not done XD

OpenStudy (anonymous):

don't worry, it's a.m. here. lol. Let's work together & we'll work through this ok?

OpenStudy (anonymous):

that's two a.m.

OpenStudy (anonymous):

so, now you understand the first one?

OpenStudy (anonymous):

sure thanks for helping ^_^ yup

OpenStudy (anonymous):

w8...it 2 am there right? sorry for bothering...if u want u can always read it tom ^^ its not due till monday :)

OpenStudy (anonymous):

Now, 2)\[\Large Ft=\frac{m(v-u)}{t}*t=m(v-u)\] You know the force, the time, the mass, the initial velocity. Plug them in to get the final velocity.

OpenStudy (anonymous):

No problem, I am sort of a night person.

OpenStudy (anonymous):

lol, me too... except we have an appointment tom and my mom's making me sleep T.T...

OpenStudy (anonymous):

lol.

OpenStudy (anonymous):

the V is the initial velocity right?..can I ask what's the" U " is it the 250 N? that means: 500g(0m/s-250N) =~125 000? O.O

OpenStudy (anonymous):

Now, for the third. If you push the wood with that force, the net force acting on the wood will be\[\Large F_{net}=m_{wood}a=\mu_smg-\mu_kmg\\\Large\Rightarrow a=g(\mu_s-\mu_k)\].

OpenStudy (anonymous):

So, counter-intuitively, the acceleration is independent of the mass of the block!

OpenStudy (anonymous):

3.) 9.8(0.5-0.3)= 1.96

OpenStudy (anonymous):

yup. :)

OpenStudy (anonymous):

O.O...and im done with my homework thank you >.<...lol XD better get going be4 my mom unplugs the net :) thanks for ur help ^^

OpenStudy (anonymous):

Sure thing. Bye! XD

OpenStudy (anonymous):

Baybay :3 lol

OpenStudy (anonymous):

byyye...